OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 40 miles per hour (mph), and the other car is traveling 20 mph. How long will it take before the cars are 300 miles apart?

OpenStudy (anonymous):

How far will they be after one hour?

OpenStudy (anonymous):

combined rate is \(20+40=60\) mph use \[T=\frac{D}{R}\] to find the answer

OpenStudy (anonymous):

T1 = time for first car D1 = distance for first car = 40mph * T1 T2 = time for second car D2 = distance for second car = 20mph * T2 For this problem, T1 = T2 since they started at the same time. DT = sum of both distances = D1 + D2 = 300 Set up our equation. DT = 300 = D1 + D2 = (40mph * T1) + (20mph * T2) 300 = (40mph * T1) + (20mph * T2) Remembering that T1 = T2, so we can substitute either value for both values. 300 = (40mph * T1) + (20mph * T1) 300 = T1(40 + 20) Now finish it! Remember that T1 will be in hours.

OpenStudy (anonymous):

so its 300 = (5) (60). 300 =300

OpenStudy (anonymous):

Nett speed = 40+20 =60mph 300 /60 = 5 hours ANSWER

OpenStudy (anonymous):

Well that was the simple way.. Lo, @ZhangYan

OpenStudy (anonymous):

lol*

OpenStudy (anonymous):

lol im about to put another question up that i need help

OpenStudy (anonymous):

Okay!